OpenStudy (anonymous):

Solve the equation 1 over 3 x + 2 = 10 for x.

OpenStudy (anonymous):

1/3x+2=10?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x=1/24

OpenStudy (anonymous):

thank you :) very much very helpful

OpenStudy (anonymous):

but

OpenStudy (anonymous):

but what

OpenStudy (anonymous):

if 1/3 is i fraciton form and the x is by the 3 then it is x=1/24

OpenStudy (anonymous):

if the x is after the 1/3 equation then the answer is x - 24

OpenStudy (anonymous):

yes so its just 24

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

so its 1 over 3, space X, the x would be in the midle where the mid line is correct

OpenStudy (anonymous):

if its in the midline then its x=24

OpenStudy (anonymous):

yes your correct